OpenStudy (anonymous):

Check this.? Im not very good with logarithms and i watched a video and it said to do it like this but im not sure Log\/2(3)+log\/2(x)=3 then cancel out the logs as they are the same base so, 3+x=3 -3 -3 x=0???

Nnesha (nnesha):

what's the original question ?

OpenStudy (anonymous):

Log\/2(3)+log\/2(x)=3

Nnesha (nnesha):

i mean equation \[\log_2(3)+\log_2 x=3\] like this OR

Nnesha (nnesha):

\[\frac{ \log_2 3 }{ \log_2 x} =3\]

Nnesha (nnesha):

maybe or maybe division hmm which one is correct ? graysongraddyLOL

OpenStudy (anonymous):

the first one is the original

Nnesha (nnesha):

oh okay familiar with the log properties ??

OpenStudy (anonymous):

Not exactly.

Nnesha (nnesha):

quotient rule\[\large\rm log_b x - \log_b y = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\large\rm log_b x^y = y \log_b x\]

Nnesha (nnesha):

look at that and tell me which one you should apply ?

OpenStudy (anonymous):

The last one.?

OpenStudy (anonymous):

Math just absolutely dumbfounds me.

Nnesha (nnesha):

hmm alright there is PLUS sign between log now look at the log identities

OpenStudy (anonymous):

the subexponent of 2 is the identity correct.?

Nnesha (nnesha):

2 is base

Nnesha (nnesha):

quotient rule\[\large\rm\color{reD}{ log_b x - \log_b y = \log_b \frac{ x }{ y}}\] to condense you can change subtraction to division product rule \[\large\rm \color{reD}{log_b x + \log_b y = \log_b( x \times y )}\] addition ----> multiplication power rule \[\large\rm \color{Red}{log_b x^y = y \log_b x}\] you don't need ^power rule for this e quation we should apply it when there is a number at front of log

OpenStudy (anonymous):

so with the equation that i am using, How do i get a Log on the other side of the equal sign.?

Nnesha (nnesha):

alright we need to apply product rule which is \[\large\rm \color{reD}{log_b x + \log_b y = \log_b( x \times y )}\] log_b same base so take it out and multiply x y \[\log_b x+ \log_b y\] \[\log_b (x \times y)\]

OpenStudy (anonymous):

So multiply 3 and x.?

Nnesha (nnesha):

\[\huge\rm log_2(3)+\log_2 x=3\] like this OR yes right!!

Nnesha (nnesha):

that's how we should condense log equations \[\huge\rm log_2 (3 \times x)=3\] now we should convert log to exponential form

Nnesha (nnesha):